Terri and Juan earned $1,150 from a one-year investment of $15,000. If Terri invested some of the money at 8% interest and Juan invested the rest at 7%, how much did each invest?
Answer by mananth(16949)

Terri 8.00%per annum ------------- Amount invested =x
Juan 7.00%per annum ------------ Amount invested = y
15000
Interest----- 1150
Sum of investments
x + y= 15000 -------------1
Interest earned on both investments
8.00% x + 7.00% y= 1150
Multiply by 100
8 x + 7 y= 115000.00 --------2
Multiply (1) by -8
we get
-8 x -8 y= -120000.00
Add this to (2)
0 x -1 y= -5000
divide by -1
y = 5000
Terri 7.00% $ 5000
Juan 8.00% $ 10000

CHECK
5000 --------- 7.00% ------- 350.00
10000 ------------- 8.00% ------- 800.00
Total -------------------- 1150.00

m.ananth@hotmail.ca